OpenStudy (anonymous):

If f(x)=2x-6 f(2) can someone tell me how to solve it! please

OpenStudy (anonymous):

where you see an \(x\) put a \(2\) and compute \[f(2)=2\times 2-6\]

OpenStudy (anonymous):

that's it!

OpenStudy (anonymous):

yeh and remember multiply first :D

OpenStudy (anonymous):

im lost right now . >_<

OpenStudy (anonymous):

yeah that's it!

OpenStudy (anonymous):

if you can compute \[2\times 2-6\] then you can do this

OpenStudy (anonymous):

your looking for f(2), so wherever x is replace it with 2 and u get f(2)=(2×2)−6

OpenStudy (anonymous):

oh ok maybe i got it thank u so much

OpenStudy (anonymous):

and شكرا

OpenStudy (anonymous):

كنت موضع ترحيب

OpenStudy (anonymous):

i can't see all of that

OpenStudy (anonymous):

al'afw

OpenStudy (anonymous):

wow both of u arabic cool lol

OpenStudy (anonymous):

where u from satellite73

OpenStudy (anonymous):

where r u from magepker728

OpenStudy (anonymous):

Libya

OpenStudy (anonymous):

u?

OpenStudy (anonymous):

ahlan Iraq

OpenStudy (anonymous):

kewl tasharafna

OpenStudy (xapproachesinfinity):

Hola hola! @satellite73 doesn't speak Arabic he is just cheating hehehe

OpenStudy (anonymous):

man 3nta @xapproachesinfinity

OpenStudy (anonymous):

@magepker728 can u help me again please!

OpenStudy (xapproachesinfinity):

Ana insan hehe

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

If f(x)=2x-6 f(-1/2)

OpenStudy (anonymous):

2(-1/2)-6 let me know if u can solve it

OpenStudy (anonymous):

=-7? right

OpenStudy (xapproachesinfinity):

Good!

OpenStudy (anonymous):

u got it :D way to go

OpenStudy (anonymous):

yay thx so much

OpenStudy (anonymous):

anytime

OpenStudy (xapproachesinfinity):

f(x)=x^3-x^2+3 what is f(3)

OpenStudy (anonymous):

ur 69 u can solve it :D

OpenStudy (anonymous):

min ayna 3nta @xapproachesinfinity ?

OpenStudy (xapproachesinfinity):

i can't solve it^_^ 69 means nothing

OpenStudy (xapproachesinfinity):

من المغرب حبيبي^_^

OpenStudy (anonymous):

@magepker728 can u help me again please!

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

g(x)=x-2x^2 g(3)

OpenStudy (anonymous):

oooh maghrebiya tu parle le francais donc non?

OpenStudy (xapproachesinfinity):

Mina elmaghrib! this open study stops from working sometimes lol

OpenStudy (anonymous):

3-2(9)

OpenStudy (xapproachesinfinity):

Oh non je parle pas la langue francais!

OpenStudy (anonymous):

lol et l'espagnole?

OpenStudy (anonymous):

@Tuta u got it?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

3-(2x9)

OpenStudy (xapproachesinfinity):

meme! je parle justement la langue anglais, arabe, amazighia

OpenStudy (anonymous):

i didn't understand from where u got the nine

OpenStudy (xapproachesinfinity):

he/she square the 3 it became 9

OpenStudy (xapproachesinfinity):

you have x-2x^2 remember

OpenStudy (anonymous):

that is because wherever x is u replace it with 3 and the x^2=3^2

OpenStudy (anonymous):

oh ok thank u

OpenStudy (anonymous):

anytime

OpenStudy (anonymous):

@xapproachesinfinity i speak a little bit of tamasheq

OpenStudy (xapproachesinfinity):

Oh yeah! imprissive! so you speak {English, French, Arabic, Spanish...}

OpenStudy (anonymous):

yes sir :D been traveling around since i was little lol

OpenStudy (xapproachesinfinity):

You are quite linguistic are you? that's amusing! my french sucks! i completely forgot what i know hehe

OpenStudy (anonymous):

i have been to morocco once.. nice country to be honest, was there for 3 weeks and let me tell you one of the best experiences i had ever had

OpenStudy (xapproachesinfinity):

Oh yeah! where did you go?

OpenStudy (anonymous):

we stayed in marrakesh

OpenStudy (anonymous):

then to cassablanca which i found really amazing just the name says lol most houses over there were colored white lol

OpenStudy (anonymous):

interrupt your conversation but @magepker728 i need your help again

OpenStudy (anonymous):

sure

OpenStudy (xapproachesinfinity):

hehehe, Didn't you find Marrakesh as well all red! anyways i don't like Casablanca I like Tangier, Araich... Good to know you "Hassan" is my name! i will see you later i have to go now

OpenStudy (anonymous):

alright brother name and the name is Abdoulkarim

OpenStudy (anonymous):

changed each equation to function notation and then find the value equation function notation find value for... y=3x^2-2 f(2)

OpenStudy (xapproachesinfinity):

pleasure meeting ya see ya around

OpenStudy (anonymous):

so the notation is going to be f(2)=3x^2-2 (3x4)-2=?

OpenStudy (anonymous):

can u do it step by step so i'll understand it better and understand how to work the others please!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since the equation is already already in its formal of y=....., all you have to is replace y= to f(x), which will give you f(x)=3x^2-2 and wherever there is an x replace it with 2 since your looking for f(2)

OpenStudy (anonymous):

then youll get f(2)=3(2)^2-2 fallow the rule of pemdas to solve in this case exponent first, so (3x4) and then -2

OpenStudy (anonymous):

oh ok so is that equal 10 or im lost?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

oh ok thank u im just so bad at math that's why i keep asking sorry for taking ur time

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

can you help me again i know I'm asking too much :( @magepker728

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

y=x/3+1 and find the value for f(6) i was working in on it i guess it's equal 3! is that right

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

i mean i was working on it lol

OpenStudy (anonymous):

3 isnt the answer :D

OpenStudy (anonymous):

uh how :((

OpenStudy (anonymous):